an inclined plane making an angle of 25 degree with the horizontal has a pulley at its top. A 30kg block on plane is connected to a freely hanging 20kg block by means of a string passing over the pulley. Compute the distance the 20kg block will fall in 2 second starting from rest. Neglect the friction.


1
Expert's answer
2021-10-18T11:08:22-0400

Let

M"_A" = 30 kg

M"_B" = 20 kg


For Block B

"M_Bg-T=M_ba"

"20\u00d79.8-T=20a"

"T+20a=196.......i)"


For Block A

"T-M_Agsin25=M_aa"

"T-30\u00d79.8\u00d7sin25=30a"a

"T-30a=124.25......ii)"

"i-ii"

"50a=196-124.25"

"a=1.435m\/s^2"

So distance traveled by 20 kg block is

"Y=V_{oy}t+\frac{1}{2}at^2=0+\frac{1}{2}×1.435×2^2"

"Y=2.87m"